# Entanglement assisted probe of the non-Markovian to Markovian transition in open quantum system dynamics

### v1

@@ Abstract
We utilize a superconducting qubit processor to experimentally probe the transition from non-Markovian to Markovian dynamics of an entangled qubit pair. We prepare an entangled state between two qubits and monitor the evolution of entanglement over time as one of the qubits interacts with a small quantum environment consisting of an auxiliary transmon qubit coupled to its readout cavity. We observe the collapse and revival of the entanglement as a signature of quantum memory effects in the environment. We then engineer the non-Markovianity of the environment by populating its readout cavity with thermal photons to show a transition from non-Markovian to Markovian dynamics, reaching a regime where the quantum Zeno effect creates a decoherence-free subspace that effectively stabilizes the entanglement between the qubits.
@@
Non-Markovian dynamics have the potential to enable novel applications stemming from memory effects in the environment, such as new approaches towards fault-tolerant quantum computation [16, 17, 18], fidelity improvement in the implementation of the teleportation algorithms [19], and coherence preservation [20]. The non-Markovianity of an open quantum system can be measured using two common methods. The most prominent measure is known as the trace distance method proposed by Breuer et. al. [21], which relies on the fact that any completely positive trace-preserving (CPTP) quantum map between two-time steps will only result in a decrease of the distinguishability between two quantum states, hence any increase in the distance between states is associated with memory effects [22]. Later, an entanglement measure was introduced by Rivas et. al. [23], where one probes quantum memory effects by allowing part of an entangled pair to interact with an environment. Again, a CPTP map will only decrease the degree of entanglement and an increase in entanglement during the system evolution is a signature of quantum memory effects. Both methods [24] have been employed to observe signatures of non-Markovianity, notably in nitrogen-vacancy centers [25, 26, 27], photonic systems [28, 29, 30], and the trapped ion platform [31].
In this Letter, we harness the entanglement between two superconducting qubits as a probe of quantum memory effects. We initialize the qubits in a Bell state and study the qubits’ concurrence [32] over time as one of the qubits interacts with a small quantum environment consisting of a third qubit dispersively coupled to a microwave resonator. We observe collapse and then revival of the qubits’ concurrence as a clear signature of the non-Markovian nature of the environment as the qubit becomes entangled and then disentangled with the environment. The non-Markovianity of the environment is then tuned by introducing Lindblad dephasing on the environment. This allows us to investigate a non-Markovian–to–Markovian transition in the dynamics. In this non-Markovian regime, we further increase the dissipation on the environment, ultimately reaching a regime where the quantum Zeno effect pins the environment state, thereby preserving the qubits’ entanglement.
@@
Figure 1(a) displays the basic setup of the experiment. The experiment comprises a three-qubit processor with individual readout resonators dispersively coupled to each qubit and nearest-neighbor qubits sharing a resonator mediated coupling. The readout resonators allow us to perform individual state readouts of the three qubits by probing the associated microwave resonators with a microwave drive. We first focus on a sub-portion of the processor with two qubits denoted as the “Qubit” and the “Ancilla”. The Qubit is frequency tunable via a Superconducting QUantum Interference Device (SQUID) loop and the Ancilla is fixed-frequency, both designed to be in the transmon regime [33]. In order to minimize the decoherence effects from flux noise, we operate the Qubit at its flux sweet spot and introduce coupling to the Ancilla via parametric modulation [34]. To this end, we apply an ac radio frequency drive on the Qubit fast flux line at roughly half the detuning between the Qubit and Ancilla (Fig. 1(b)). We identify the resonance condition between the Qubit and Ancilla by initializing the qubit in its excited state and then applying the parametric modulation for a variable duration. Figure 1(c) shows the time evolution of the Ancilla $\langle\sigma_{z}\rangle\equiv Z_{A}$ near the parametric resonance. We observe a clear chevron profile with detuning from which we extract a parametric coupling rate of $\Omega_{\mathrm{Q,A}}/2\pi=0.477$ MHz.
@@
With the Qubit and Ancilla entangled, we now study the evolution of the entanglement over time as the system sits idle. We display the Qubit–Ancilla concurrence versus time in Fig. 2(c). The concurrence slowly decreases over a timescale consistent with the respective individual dephasing times of the Qubit ($T_{2}^{*}=39\ \mu$s) and Ancilla ($T_{2}^{*}=41\ \mu$s), as given by the dashed line in Fig. 2(c).
@@
Next, we investigate the time evolution of the Qubit–Ancilla concurrence for different values of the Environment dephasing. Increasing the Environment dephasing induces a transition from non-Markovian to Markovian dynamics as displayed in Fig. 4(d). We quantify this transition via the measure (1) as displayed in Fig. 4(e); as the dephasing of the Environment is increased beyond $\gamma\simeq 1\ (\mu\mathrm{s})^{-1}$ the dynamics are Markovian.
@@
In Fig. 4(f) we display the concurrence versus time for a few selected values of $\gamma$. We note two important trends; first, in the non-Markovian regime, increasing dephasing accelerates the decay envelope of the concurrence (compare $\gamma=0$ and $\gamma=0.5$), and second, in the Markovian regime, further increasing dephasing slows the decay of the concurrence ($\gamma=2.4$ and $\gamma=6.4$). This can be understood within the context of the quantum Zeno effect [41, 42, 43, 44, 45, 45, 46, 47]. The thermal photons perform measurement (at rate $\gamma$) of the Environment, which slows the coupling induced by the parametric drive. In Fig. 5, we explore in detail how the dephasing of the Environment affects the Qubit–Ancilla entanglement. Figure 5a displays the concurrence versus time for several values of the dephasing in the Markovian regime. For increasing measurement on the Environment, we see that the decay of concurrence is slowed, approaching the limiting case where the Qubit is completely uncoupled from the Environment. We characterize the exponential decay of the concurrence with a rate $\Gamma_{\mathrm{c}}$, and display this rate versus Environment dephasing in Fig. 5b in both the non-Markovian and Markovian regimes; the transition between these two regimes coincides with the onset of Zeno stabilization of the entanglement. Under the standard analysis of the Zeno effect [48, 46], we expect $\Gamma_{\mathrm{c}}=\Omega_{\mathrm{Q,E}}^{2}/4\gamma+\Gamma_{0}$. Here $\Gamma_{0}=1/T_{2}^{*Q}+1/T_{2}^{*A}$ is the decay rate of the concurrence when the Environment is decoupled. We observe close agreement with this expected scaling (red curve). This demonstrates a new approach to preserving quantum entanglement via Zeno-enabled pinning of environment states.
In conclusion, we have quantified the non-Markovian to Markovian transition with an entanglement-assisted probe. Importantly, the probe is sensitive to the quantum memory of the environment; a classical environment that stores populations will not result in the revival of concurrence for the entangled probe. This approach can have utility in the test of the quantum nature of decoherence channels (e.g. in testing models of quantum gravity [49]). Moreover, by introducing controllable dissipation on the environment we observe stabilization of the Qubit–Ancilla subsystem, highlighting how dissipation forms a powerful tool for quantum subspace engineering [13].
@@ Acknowledgements.
Acknowledgments—The authors are thankful for the useful discussions with Kade Head-Marsden, Patrick Harrington, Weijian Chen, Kaiwen Zheng, Maryam Abbasi, Serra Erdamar, and Archana Kamal. This research was supported by NSF Grants No. PHY-1752844 (CAREER) and OMA-1936388, the Air Force Office of Scientific Research (AFOSR) Multidisciplinary University Research Initiative (MURI) Award on Programmable systems with non-Hermitian quantum dynamics (Grant No. FA9550-21-1- 0202), the John Templeton Foundation, Grant No. 61835, ONR Grant Nos. N00014-21-1-2630 and N00014-21-1-2688 (YIP), Research Corp. Grant No. 27550 (Cottrell), European Union’s FET Open AVaQus grant no. 899561 and Marie Sklodowska-Curie grant no. MSCA-IF-835791. Devices were fabricated and provided by the Superconducting Qubits at Lincoln Laboratory (SQUILL) Foundry at MIT Lincoln Laboratory, with funding from the Laboratory for Physical Sciences (LPS) Qubit Collaboratory.
@@
I Calibration of the Qubit-Environment parametric coupling
@@ II Experimental Setup
III Device design and simulations
@@
IV Two-qubit state tomography

### v2

# thanks: Equal contributionthanks: Equal contribution

@@ Abstract
We utilize a superconducting qubit processor to experimentally probe non-Markovian dynamics of an entangled qubit pair. We prepare an entangled state between two qubits and monitor the evolution of entanglement over time as one of the qubits interacts with a small quantum environment consisting of an auxiliary transmon qubit coupled to its readout cavity. We observe the collapse and revival of the entanglement as a signature of quantum memory effects in the environment. We then engineer the non-Markovianity of the environment by populating its readout cavity with thermal photons to show a transition from non-Markovian to Markovian dynamics, ultimately reaching a regime where the quantum Zeno effect creates a decoherence-free subspace that effectively stabilizes the entanglement between the qubits.
@@
Non-Markovian dynamics have the potential to enable novel applications stemming from memory effects in the environment, such as new approaches towards fault-tolerant quantum computation [16, 17, 18], quantum control [19], fidelity improvement in the implementation of the teleportation algorithms [20], and coherence preservation [21]. The non-Markovianity of an open quantum system can be measured using two common methods [22]. The most prominent measure is known as the trace distance method proposed by Breuer et. al. [23], which relies on the fact that any completely positive trace-preserving (CPTP) quantum map between two-time steps will only result in a decrease of the distinguishability between two quantum states, hence any increase in the distance between states is associated with memory effects [24]. Later, an entanglement measure was introduced by Rivas et. al. [25], where one probes quantum memory effects by allowing part of an entangled pair to interact with an environment. Again, a CPTP map will only decrease the degree of entanglement and an increase in entanglement during the system evolution is a signature of quantum memory effects. Both methods [26] have been employed to observe signatures of non-Markovianity, notably in nitrogen-vacancy centers [27, 28, 29, 30], photonic systems [31, 32, 33, 34, 35], nuclear magnetic resonance [36, 35, 37], trapped ions [38], and on superconducting processors [39].
In this Letter, we harness the entanglement between two superconducting qubits as a probe of quantum memory effects. We initialize the qubits in a Bell state and study the qubits’ concurrence [40] over time as one of the qubits interacts with a small quantum environment consisting of a third qubit dispersively coupled to a microwave resonator. We observe collapse and then revival of the qubits’ concurrence as a clear signature of the non-Markovian nature of the environment as the qubit becomes entangled and then disentangled with the environment. The non-Markovianity of the environment is then tuned by introducing Lindblad dephasing on the environment [41]. This allows us to investigate a transition away from non-Markovian dynamics to a regime where the GKSL master equation describes the dynamics. Since the GKSL master equation requires a Markovian approximation, we refer to this regime as the Markovian regime. In this Markovian regime, we further increase the dissipation on the environment, ultimately reaching a regime where the quantum Zeno effect pins the environment state, thereby preserving the qubits’ entanglement.
@@
Figure 1(a) displays the basic setup of the experiment, with the system Hamiltonian given in [42]. The experiment comprises a three-qubit processor with individual readout resonators dispersively coupled to each qubit and nearest-neighbor qubits sharing a resonator mediated coupling. The readout resonators allow us to perform individual state readouts of the three qubits by probing the associated microwave resonators with a microwave drive. We first focus on a sub-portion of the processor with two qubits denoted as the “Qubit” and the “Ancilla”. The Qubit is frequency tunable via a Superconducting QUantum Interference Device (SQUID) loop and the Ancilla is fixed-frequency, both designed to be in the transmon regime [43]. In order to minimize the decoherence effects from flux noise, we operate the Qubit at its flux sweet spot and introduce coupling to the Ancilla via parametric modulation [44]. To this end, we apply an ac radio frequency drive on the Qubit fast flux line at roughly half the detuning between the Qubit and Ancilla (Fig. 1(b)). We identify the resonance condition between the Qubit and Ancilla by initializing the qubit in its excited state and then applying the parametric modulation for a variable duration. Figure 1(c) shows the time evolution of the Ancilla $\langle\sigma_{z}\rangle\equiv Z_{A}$ near the parametric resonance. We observe a clear chevron profile with detuning from which we extract a parametric coupling rate of $\Omega_{\mathrm{Q,A}}/2\pi=0.477$ MHz.
@@
With the Qubit and Ancilla entangled, we now study the evolution of the entanglement over time as the system sits idle. We display the Qubit–Ancilla concurrence versus time in Fig. 2(c). The concurrence slowly decreases over a timescale consistent with the respective individual dephasing times of the Qubit ($T_{2}^{*(Q)}=39\ \mu$s) and Ancilla ($T_{2}^{*(A)}=41\ \mu$s), e.g. $\mathcal{C}\propto\exp(-t/T_{2}^{*(Q)}-t/T_{2}^{*(A)})$, as given by the dashed line in Fig. 2(c).
@@
Next, we investigate the time evolution of the Qubit–Ancilla concurrence for different values of the Environment dephasing. Increasing the Environment dephasing induces a transition from non-Markovian to Markovian dynamics as displayed in Fig. 4(d). We quantify the transition away from non-Markovian dynamics via the measure (Eq. 1) as displayed in Fig. 4(e); as the dephasing of the Environment is increased beyond $\gamma\simeq 1\ (\mu\mathrm{s})^{-1}$, $\mathcal{N}$ becomes consistent with zero. However, the dynamics are not immediately Markovian, which we define by the applicability of the GKSL master equation to the Qubit–Ancilla subsystem. As we study in [42], the GKSL master equation yields exponentially decaying dynamics of the concurrence. This matches well the measured dynamics for $\gamma\gtrsim 3\ (\mu\mathrm{s})^{-1}$, but fails to capture the dynamics for smaller values of $\gamma$. As such, the transition between these two regions, as defined, is not abrupt.
@@
In Fig. 4(f) we display the concurrence versus time for a few selected values of $\gamma$. We note two important trends; first, in the non-Markovian regime, increasing dephasing accelerates the decay envelope of the concurrence (compare $\gamma=0$ and $\gamma=0.5$), and second, in the Markovian regime, further increasing dephasing slows the decay of the concurrence ($\gamma=2.4$ and $\gamma=6.4$). This can be understood within the context of the quantum Zeno effect [50, 51, 52, 53, 54, 54, 55, 56, 57]. The thermal photons perform measurement (at rate $\gamma$) of the Environment, which slows the coupling induced by the parametric drive. In Fig. 5, we explore in detail how the dephasing of the Environment affects the Qubit–Ancilla entanglement. Figure 5a displays the concurrence versus time for several values of the dephasing in the Markovian regime. For increasing measurement on the Environment, we see that the decay of concurrence is slowed, approaching the limiting case where the Qubit is completely uncoupled from the Environment. We characterize the exponential decay of the concurrence with a rate $\Gamma_{\mathrm{c}}$, and display this rate versus Environment dephasing in Fig. 5b in both the non-Markovian and Markovian regimes; the transition between these two regimes coincides with the onset of Zeno stabilization of the entanglement. Under the standard analysis of the Zeno effect [58, 55], we expect $\Gamma_{\mathrm{c}}=\Omega_{\mathrm{Q,E}}^{2}/4\gamma+\Gamma_{0}$. Here $\Gamma_{0}=1/T_{2}^{*Q}+1/T_{2}^{*A}$ is the decay rate of the concurrence when the Environment is decoupled. We observe close agreement with this expected scaling (red curve). This demonstrates a new approach to preserving quantum entanglement via Zeno-enabled pinning of environment states.
In conclusion, we have quantified the transition from non-Markovian dynamics to Zeno dynamics with an entanglement-assisted probe. Importantly, the probe is sensitive to the quantum memory of the environment; a classical environment that stores populations will not result in the revival of concurrence for the entangled probe. This approach can have utility in the test of the quantum nature of decoherence channels (e.g. in testing models of quantum gravity [59]). Moreover, by introducing controllable dissipation on the environment we observe stabilization of the Qubit–Ancilla subsystem, highlighting how dissipation forms a powerful tool for quantum subspace engineering [13].
@@ Acknowledgements.
Acknowledgments—The authors are thankful for the useful discussions with Daniel Lidar, Kade Head-Marsden, Patrick Harrington, Weijian Chen, Kaiwen Zheng, Maryam Abbasi, Serra Erdamar, and Archana Kamal. This research was supported by NSF Grants No. PHY-1752844 (CAREER) and OMA-1936388, the Air Force Office of Scientific Research (AFOSR) Multidisciplinary University Research Initiative (MURI) Award on Programmable systems with non-Hermitian quantum dynamics (Grant No. FA9550-21-1- 0202), the John Templeton Foundation, Grant No. 61835, ONR Grant Nos. N00014-21-1-2630 and N00014-21-1-2688 (YIP), Research Corp. Grant No. 27550 (Cottrell), European Union’s FET Open AVaQus grant no. 899561 and Marie Sklodowska-Curie grant no. MSCA-IF-835791. Devices were fabricated and provided by the Superconducting Qubits at Lincoln Laboratory (SQUILL) Foundry at MIT Lincoln Laboratory, with funding from the Laboratory for Physical Sciences (LPS) Qubit Collaboratory.
@@
I System Hamiltonian
The Hamiltonian of the three qubit system can be expressed in terms of the Pauli operators $\sigma_{x}$ and $\sigma_{z}$ for the three qubits as ($\hbar=1$),
$H=\sum_{i=\mathrm{A,Q,E}}\big{(}-\frac{1}{2}\tilde{\omega}_{\mathrm{q},i}\sigma_{z}^{i}+(\omega_{\mathrm{c},i}-\chi_{\mathrm{qc},i})a^{\dagger}_{i}a_{i}\sigma_{z}^{i}\big{)}+J_{\mathrm{A,Q}}\sigma_{x}^{\mathrm{A}}\sigma_{x}^{\mathrm{Q}}+J_{\mathrm{Q,E}}\sigma_{x}^{Q}\sigma_{x}^{E}.$ (2)
The cavity creation(annihilation) operators are denoted as $a^{\dagger}(a)$ with $\tilde{\omega}_{\mathrm{q}}=\omega_{\mathrm{q}}+\chi_{\mathrm{qc}}$ representing the Lamb-shifted qubit frequency in the dispersive frame. The dispersive coupling rates $\chi_{\mathrm{qc}}$, qubit frequencies, and cavity frequencies are given in Table 1. The coupling rates $J_{\mathrm{A,Q}}$ and $J_{\mathrm{Q,E}}$ are tuned by application of parametric modulation. In addition, the induced dephasing in the environment can be described by adding a simple Lindbladian term to the master equation governing the system’s time evolution, which can be formulated as,
$\dot{\rho}(t)=-i[H,\rho(t)]+\gamma\mathcal{L}_{\mathrm{E}}\rho(t),$ (3)
where the Lindblad operator is defined as, $\mathcal{L}_{\mathrm{E}}\rho=(\sigma_{z}^{\mathrm{E}}\rho\sigma_{z}^{\mathrm{E}}-\rho)$, $\gamma$ denotes the induced dephasing rate in the environment, and $\rho(t)$ is the system’s density matrix at time $t$.
II Numerical simulation of the Qubit–Ancilla dynamics in the Markovian regime
The Hamiltonian of the Qubit–Ancilla system in the Markovian regime can be written as ($\hbar=1$),
$H_{\mathrm{A},\mathrm{Q}}=-\frac{1}{2}\omega_{\mathrm{A}}\sigma_{z}^{\mathrm{A}}-\frac{1}{2}\omega_{\mathrm{Q}}\sigma_{z}^{\mathrm{Q}}.$ (4)
This Hamiltonian does not include any coupling between the Qubit and Ancilla, and is therefore applicable after the two qubits have been initialized in an entangled state. The dynamics of the system in presence of dephasing on both qubits is governed by a simple GKSL master equation as,
$\dot{\rho}(t)=-i[H_{\mathrm{A,Q}},\rho(t)]+\sum_{j=\mathrm{A},\mathrm{Q}}\frac{1}{2}\left(2C_{j}\rho(t)C_{j}^{\dagger}-\rho(t)C_{j}^{\dagger}C_{j}-C_{j}^{\dagger}C_{j}\rho(t)\right),$ (5)
where $C_{j}=\sqrt{\gamma_{j}}\sigma_{z}^{j}$ are the collapse operators taking into account the dephasing on each qubit with rate $\gamma_{j}$.
We study the dynamics of this simple two–qubit model and its applicability to the observed dynamics in the Markovian regime. We use the built-in mesolve function in the QuTiP package [60, 61] to numerically solve for Eq. 5 in the rotating frame to resolve the system’s density matrix at various time steps. We initialize the two-qubit $\rho$ in an entangled state consistent with the experiment and set $\gamma_{A}=0.024\ \mu\mathrm{s}^{-1}$, consistent with the Ancilla’s dephasing time. We then find the best-fit simulation by tuning the value of $\gamma_{Q}$, modeling the coupling to the environment as a simple Markovian bath. Figure 6 compares these simulation results to the experimental data presented in Fig. 5. These simulations match the experimental data for all cases except $\gamma=1.6\ \mu\mathrm{s}^{-1}$, where slight deviations from exponential decay are observed in the data. This analysis reveals a level of ambiguity about the sharpness of the transition between non-Markovian and Markovian dynamics. Indeed there exists a parameter range (in our case between $\gamma\approx 1$ and $\gamma\approx 3\ \mu\mathrm{s}^{-1}$) where the non-Markovian measure $\mathcal{N}=0$, but the GKSL master equation does not exactly match the experimental data. Detailed analysis of this transition region will be the subject of future investigation.
Figure 6: Concurrence evolution of the Qubit–Ancilla system in the Markovian regime. We compare data for the concurrence to results from the simulation of Eq. 5 for the four cases displayed in Fig. 5.
III Calibration of the Qubit-Environment parametric coupling
@@ IV Experimental Setup
V Device design and simulations
@@
VI Two-qubit state tomography